# Simple Drop-in LoRA Conditioning on Attention Layers Will Improve Your Diffusion Model

### v1

@@ 5.6 Extrapolating conditioning information
First, we input the linear interpolation $\alpha c_{i}+(1-\alpha)c_{j}$ ($0\leq\alpha\leq 1$) of two class inputs $c_{i}$ and $c_{j}$ (corresponding to ‘airplane’ and ‘horse’, respectively) to observe the continuous transition between classes. As shown in the top of Figure 5, both the scale-and-shift EDM and ClassLoRA EDM models effectively interpolate semantic information across different classes. However, when a scaled input $\beta c_{i}$ is received, with $\beta$ ranging from -1 to 1, scale-and-shift EDM generates unrecognizable images when $\beta<0$, while ClassLoRA EDM generates plausible images throughout the whole range, as shown in the bottom of Figure 5. This toy experiment shows that LoRA-based conditioning may be more robust to extrapolating conditioning information beyond the range encountered during training. Appendix C provides further details.
@@ A.3.1 CIFAR-10
For CIFAR-10 training, we use the DDPM++ Song et al. (2021b) with channel per resolution as 2, 2, 2. We use Adam optmizer with learning rate of $1\times 10^{-3}$, momentum of $(\beta_{1},\beta_{2})=(0.9,0.999)$, no weight decay and dropout rate of 0.13. We train the model with EDM preconditioning and EDM loss proposed in the paper for 200M training images (counting repetition) with batch size of 512 and augmentation probability of 0.13. For sampling, we used checkpoints saved every 50 iterations with exponential moving average of $0.99929$, and sample image using Heun’s method for 18 steps (35 NFE).
@@ A.3.2 FFHQ64
For FFHQ training, we use the DDPM++ Song et al. (2021b) with channel per resolution as 1, 2, 2, 2. We use Adam optmizer with learning rate of $2\times 10^{-4}$, momentum of $(\beta_{1},\beta_{2})=(0.9,0.999)$, no weight decay and dropout rate of 0.05. We train the model with EDM preconditioning and EDM loss proposed in the paper for 200M training images (counting repetition) with batch size of 256 and augmentation probability of 0.15. For sampling, we used checkpoints saved every 50 iterations with exponential moving average of about 0.99965, and sample image using Heun’s method for 40 steps (79 NFE).
@@ A.5 MLP for composition weights of UC-LoRA
Appendix B Cosine similarity between $\omega_{t}$s in nano diffusion
@@
Appendix C Extrapolating conditioning information with Class LoRA
@@
Appendix D Image generation samples
D.1 IDDPM
D.1.1 Unconditional CIFAR-10
@@
D.1.2 Class-conditional CIFAR-10
@@ D.2.1 Unconditional CIFAR-10
D.2.2 Class-conditional CIFAR-10

### v3

@@ 5.6 Extrapolating conditioning information
First, we input the linear interpolation $\alpha c_{i}+(1-\alpha)c_{j}$ ($0\leq\alpha\leq 1$) of two class inputs $c_{i}$ and $c_{j}$ (corresponding to ‘airplane’ and ‘horse’, respectively) to observe the continuous transition between classes. As shown in the top of Figure 5, both the scale-and-shift EDM and ClassLoRA EDM models effectively interpolate semantic information across different classes. However, when a scaled input $\beta c_{i}$ is received, with $\beta$ ranging from -1 to 1, scale-and-shift EDM generates unrecognizable images when $\beta<0$, while ClassLoRA EDM generates plausible images throughout the whole range, as shown in the bottom of Figure 5. This toy experiment shows that LoRA-based conditioning may be more robust to extrapolating conditioning information beyond the range encountered during training. Appendix D provides further details.
@@ 6 Conclusion
7 Acknowledgements
This research was supported by a grant from KRAFTON AI. AN was supported in part by the Ministry of Science and ICT (MSIT), South Korea, under the Information Technology Research Center (ITRC) Support Program #IITP-2024-RS-2022-00156295. EKR was supported by the National Research Foundation of Korea(NRF) grant funded by the Korean government (No.RS-2024-00421203). We thank Sehyun Kwon and Dongwhan Rho for their careful reviews and valuable feedback. We also thank the anonymous reviewers for their insightful comments.
@@ A.3.1 CIFAR-10
For CIFAR-10 training, we use the DDPM++ (Song et al., 2021b) with channel per resolution as 2, 2, 2. We use Adam optmizer with learning rate of $1\times 10^{-3}$, momentum of $(\beta_{1},\beta_{2})=(0.9,0.999)$, no weight decay and dropout rate of 0.13. We train the model with EDM preconditioning and EDM loss proposed in the paper for 200M training images (counting repetition) with batch size of 512 and augmentation probability of 0.13. For sampling, we used checkpoints saved every 50 iterations with exponential moving average of $0.99929$, and sample image using Heun’s method for 18 steps (35 NFE).
@@ A.3.2 FFHQ64
For FFHQ training, we use the DDPM++ (Song et al., 2021b) with channel per resolution as 1, 2, 2, 2. We use Adam optmizer with learning rate of $2\times 10^{-4}$, momentum of $(\beta_{1},\beta_{2})=(0.9,0.999)$, no weight decay and dropout rate of 0.05. We train the model with EDM preconditioning and EDM loss proposed in the paper for 200M training images (counting repetition) with batch size of 256 and augmentation probability of 0.15. For sampling, we used checkpoints saved every 50 iterations with exponential moving average of about 0.99965, and sample image using Heun’s method for 40 steps (79 NFE).
@@ A.5 MLP for composition weights of UC-LoRA
Appendix B Compositional ClassLoRA
To verify the effectiveness of a compositional version of ClassLoRA, we conducted a proof-of-concept experiment with EDM on CIFAR-10. We finetuned a pretrained unconditional EDM for class-conditional CIFAR-10 sampling using a single LoRA basis and randomly initialized compositional weights. Note in the non-compositional ClassLoRA, we used 10 LoRA bases, one for each class. As a result, with the addition of only 61560 parameter counts, we were able to convert (fine-tune) an unconditional EDM to a conditional EDM and improve the FID score from 1.97 to 1.81.
Appendix C Cosine similarity between $\omega_{t}$s in nano diffusion
@@
Appendix D Extrapolating conditioning information with Class LoRA
@@
Appendix E Image generation samples
E.1 IDDPM
E.1.1 Unconditional CIFAR-10
@@
E.1.2 Class-conditional CIFAR-10
@@ E.2.1 Unconditional CIFAR-10
E.2.2 Class-conditional CIFAR-10